OpenStudy (anonymous):

How do I find the numerical value of one trigonometric function of x if sec x cot x = 4

OpenStudy (campbell_st):

well sec = 1/cos and cot = cos/sin so multiplying them means 1/cos * cos/sin = 1/sin so then you are looking at 1/sin = 4 so taking the reciprocals sin = 1/4 now you can find the 2 angles in the 1st and 2nd quadrants

OpenStudy (amigatour):

I found this on Google Image